# August 2025 IPMU25-0038

### v1

@@ Abstract
The $F(R)$ gravity models of inflation are revisited in light of the recent observations of cosmic microwave background radiation by Atacama Cosmology Telescope (ACT) and DESI Collaboration. A detailed study of the evolution equations in the Jordan frame is given and a new description of the slow-roll approximation in the $F(R)$-gravity-based models of inflation is proposed. It is found that all those models of inflation are significantly constrained by demanding a higher (than the Planck Telescope value) cosmological tilt $n_{s}$ of scalar perturbations and a positive running index $\alpha_{s}$ favored by ACT. It is not difficult to meet the ACT constraints on the scalar tilt $n_{s}$ by modifying the existing models of inflation, but simultaneously demanding a positive running $\alpha_{s}$ would rule out many of them. Using the proposed slow-roll approximation in the Jordan frame, we provide a new modification of the Starobinsky inflation model in the framework of $F(R)$ gravity, which satisfies all ACT constraints.
@@ 1 Introduction
Cosmological inflation is a paradigm that proposes the existence of a short (approximately between $10^{-36}$ s to $10^{-32}$ s) quasi-de-Sitter phase in the early Universe before the radiation-dominated era. This phase is characterized by an accelerating expansion of the Universe with a quick exit, and resolves major internal problems in the standard (Einstein-Friedmann) cosmology. Quantum fluctuations during inflation may be seeds of the large-scale structure formation in the Universe. A success of inflation in explaining the cosmic microwave background (CMB) radiation provides strong observational evidence for inflation. Nevertheless, inflation is still a small window into the early Universe and very high energy physics, while the detailed mechanism of inflation is still unknown despite many (over 300) theoretical models proposed in the literature [1].
Single-field models of inflation can provide a simple description of viable inflation in terms of a canonical scalar field $\phi$ with a potential $V(\phi)$, which is known as quintessence. The variety of single-field inflation models is thus parameterized by their potentials, whereas CMB observations still allow infinitely many possible potentials, even though the latter have the vanishing measure in the space of all potentials. Therefore, data-driven selection of inflaton potentials still needs to be complemented by fundamental theoretical principles and constraints to discriminate between the CMB-compatible potentials.
Quintessence is known to be classically equivalent to modified $F(R)$ gravity whose Lagrangian is given by a function $F(R)$ of the spacetime Ricci curvature scalar $R$, see e.g., Ref. [2]. Both descriptions are related by a Weyl transformation of spacetime metric, the quintessence picture is called the Einstein frame, and the modified gravity description is called the Jordan frame. The relation between those frames is non-trivial and may be used for imposing restrictions on inflaton potential from the gravity side. The most famous example is the Starobinsky inflationary model defined in the Jordan frame by the action [3, 4]
$S_{\rm Star.}={{M^{2}_{\rm Pl}}\over{2}}\int d^{4}x\sqrt{-g}\left(R+{{1}\over{6m^{2}}}R^{2}\right)~,$ (1)
@@
$V(\phi)={{3}\over{4}}M^{2}_{\rm Pl}m^{2}\left[1-\exp\left(-\sqrt{\hbox{\large{${{{{2}}\over{{3}}}}$}}}\phi/M_{\rm Pl}\right)\right]\,$ (2)
that nicely fits the CMB observations by Planck/BICEP [5, 6], though its origin from Eq. (1) is obscure in the Einstein frame. The Starobinsky model in the Jordan frame is not only geometrical by relying on General Relativity but also explains the origin of the approximate flatness of the potential $V(\phi)$ during inflation due to the scale invariance of the $R^{2}$-action [7], and has an attractor solution for inflation too, which sharply distinguishes the potential (2) among others.
At the same time, the Starobinsky description of inflation is expected to provide merely the leading (or dominant) terms in the gravitational effective action and in the inflaton potential, respectively, because both of them are subject to corrections from the higher-order terms that are more difficult to get. On the one hand, those corrections are important to account for more precise CMB measurements in the future. On the other hand, more experimental precision requires a more precise theoretical framework for CMB description. As a result of all that, single-field models of inflation become more and more constrained. The $F(R)$ gravity models of inflation, generalizing the Starobinsky model, were well studied in the literature, see e.g., Refs. [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. More recently, interest to those inflation models renewed due to the new data from the Atacama Cosmology Telescope (ACT) [21, 22] and DESI 2024 [23].
The single-field inflation models become even more constrained when one adds a production of primordial black holes (PBH) at smaller scales [24, 25]. The standard mechanism of PBH production is via engineering an nearly-inflection point in the potential $V$ [26, 27, 28]. However, less is known about PBH production with inflation in the Jordan frame, which may provide additional insights. Quantum loop corrections might invalidate many single-field models of inflation with PBH production [29], while loop calculations rely on a detailed shape of the potential $V$ [30] or that of the $F$-function [31, 32].
@@
Our paper is organized as follows. In Sec. 2, a detailed study of the evolution equations in a generic $F(R)$ gravity theory is given. The correspondence between the Jordan and Einstein frames is systematically described in Sec. 3. The inflation parameters are introduced and computed in Secs. 4 and 5. Also in Sec. 5 we propose new explicit formulae for the slow-roll approximation in the Jordan frame. The Starobinsky model of inflation and its predictions for CMB are reproduced in the Jordan frame in Sec. 6 from the viewpoint of our new SR approximation. The CMB measurements by Planck/BICEP Collaborations [6] and Atacama Cosmology Telescope (ACT) [21, 22] combined with DESI 2024 data [23] are discussed in Sec. 7. Our new model of Starobinsky inflation in $F(R)$ gravity is presented in Sec. 8. We summarize our conclusions in Sec. 9.
@@ 2 Evolution equations in $F(R)$ gravity models
Let us consider a generic $F(R)$ gravity model in four spacetime dimensions with an action (in the Jordan frame)
$S_{F}=\int d^{4}x\sqrt{-g}F(R),$ (3)
@@
Varying the action (3) with respect to the metric $g_{\mu\nu}$ gives the following equations of motion:
$F_{,R}R^{\mu}_{\nu}-{{1}\over{2}}F\delta^{\mu}_{\nu}-\left(g^{\mu\rho}D_{\rho}\partial_{\nu}-\delta^{\mu}_{\nu}\Box\right)F_{,R}=0\,$, (4)
@@
Equation (4) leads to the trace equation
$F_{,R}R-2F+3\Box F_{,R}=0\,.$ (5)
@@
$3\left(\dot{H}+H^{2}\right)F_{,R}-{{1}\over{2}}F=3HF_{,RR}\dot{R}\,$, (9)
while the trace equation (5) is a consequence of Eq. (9).
To describe cosmic evolution during inflation, we use the e-folds number $N=-\ln(a/a_{e})$, where $a_{e}$ is a constant. Using the relation ${{d}\over{dt}}=-H{{d}\over{dN}}$, we find that Eq. (9) is equivalent to
$\left(H^{2}\right)^{\prime}=4H^{2}-{{R}\over{3}}\,,\qquad{R\,}^{\prime}={{{}-1}\over{6H^{2}F_{,RR}}}\left[\left(R-6H^{2}\right)F_{,R}-F\right],$ (10)
@@
It is possible to rewrite Eq. (9) to the form similar to the $(0,0)$-equation in the Einstein frame, by using the procedure proposed in Ref. [34] and the following functions:
$Y={{M_{\mathrm{Pl}}}\over{\sqrt{2F_{,R}}}}\left(H+{{F_{,RR}\dot{R}}\over{2F_{,R}}}\right),$ (11)
$A={{3M_{\mathrm{Pl}}^{4}F_{,RR}^{2}}\over{4F_{,R}^{3}}}\,,\qquad V_{\mathrm{eff}}={{M_{\mathrm{Pl}}^{4}}\over{4F_{,R}^{2}}}\left(RF_{,R}-F\right)\,.$ (12)
Equation (9) implies that these functions satisfy the equation
$3M_{\mathrm{Pl}}^{2}Y^{2}={{A}\over{2}}\dot{R}^{2}+V_{\mathrm{eff}}$ (13)
@@
$\begin{split}3M_{\mathrm{Pl}}^{2}Y^{2}=&{{3M_{\mathrm{Pl}}^{4}}\over{2F_{,R}}}\left(H^{2}+{{HF_{,RR}\dot{R}}\over{F_{,R}}}+{{F_{,RR}^{2}}\over{4F^{2}_{,R}}}\dot{R}^{2}\right)\\ =&{{3M_{\mathrm{Pl}}^{4}}\over{2F_{,R}}}H^{2}+{{3M_{\mathrm{Pl}}^{4}}\over{2F_{,R}^{2}}}\left[\left(\dot{H}+H^{2}\right)F_{,R}-{{F}\over{6}}\right]+{{A}\over{2}}\dot{R}^{2}\\ =&{{A}\over{2}}\dot{R}^{2}+V_{\mathrm{eff}}\,.\end{split}$ (14)
@@ 3 Correspondence between Jordan and Einstein frames
Let us demonstrate that the function $Y(t)$ is the Hubble parameter in the Einstein frame as a function of the cosmic time in the Jordan frame. The $F(R)$ gravity action (3) can be rewritten as
$S_{J}=\int d^{4}x\sqrt{-g}\left[U(\sigma)R-V(\sigma)\right]\,$, (15)
@@
$U(\sigma)=F_{,\sigma}(\sigma),\qquad V(\sigma)=\sigma F_{,\sigma}-F\,$, (16)
and the commas denote the derivatives with respect to a given scalar field. When $F_{,\sigma\sigma}(\sigma)\neq 0$, the field $\sigma$ can be eliminated via its algebraic equation of motion, $R=\sigma$, which yields back the action (3). It follows from the definitions of $U$ and $V$ that
$V_{,\sigma}=\sigma U_{,\sigma}=\sigma F_{,\sigma\sigma}\,.$ (17)
@@
$2U\dot{H}-H\dot{U}+\ddot{U}=0.$ (19)
After the Weyl transformation of metric from the Jordan frame to the Einstein frame as
$g^{(E)}_{\mu\nu}={{2F_{,\sigma}(\sigma)}\over{M^{2}_{\mathrm{Pl}}}}g_{\mu\nu}\,$, (20)
@@
$S_{E}=\int d^{4}x\sqrt{-g^{(E)}}\left[{{M^{2}_{\mathrm{Pl}}}\over{2}}R_{E}-{{h}\over{2}}{g^{(E)\mu\nu}}\partial_{\mu}{\sigma}\partial_{\nu}{\sigma}-V_{eff}\right],$ (21)
@@
$h(\sigma)={{3M^{2}_{\mathrm{Pl}}}\over{2F_{,\sigma}^{2}}}F_{,\sigma\sigma}^{2}={{2U(\sigma)}\over{M^{2}_{\mathrm{Pl}}}}\,A(\sigma)$ (22)
@@
$V_{eff}(\sigma)={{M_{\mathrm{Pl}}^{4}}\over{4F_{,\sigma}^{2}}}\left(RF_{,\sigma}-F\right)={{M_{\mathrm{Pl}}^{4}V}\over{4U^{2}}}~.$ (23)
@@
$H_{E}={{d\ln(a_{E})}\over{dt_{E}}}={{M_{\mathrm{Pl}}}\over{\sqrt{2U}}}\left(H+{{\dot{U}}\over{2U}}\right),$ (24)
@@
${{dt_{E}}\over{dt}}={{\sqrt{2F_{,R}}}\over{M_{\mathrm{Pl}}}}\,.$ (25)
Equations (18) and (19) can be rewritten in terms of $H_{E}$ as
$3M_{\mathrm{Pl}}^{2}H_{E}^{2}={{A}\over{2}}{\dot{\sigma}}^{2}+V_{\mathrm{eff}},$ (26)
$\dot{H}_{E}={}-{{A\sqrt{2U}}\over{2M_{\mathrm{Pl}}^{3}}}\,{\dot{\sigma}}^{2},$ (27)
where the functions $A(\sigma)$ and $V_{eff}(\sigma)$ have been defined in Eq. (12).
Differentiating Eq. (26) with respect to the time $t$ and using Eq. (27) yields
$\ddot{\sigma}={}-3\sqrt{{{2U}\over{M_{\mathrm{Pl}}^{2}}}}H_{E}\dot{\sigma}-{{A_{,\sigma}}\over{2A}}{\dot{\sigma}}^{2}-{{V_{\mathrm{eff},\sigma}}\over{A}}.$ (28)
Using Eqs. (16) and (24), one gets $H_{E}=Y$, so that Eq. (26) is equivalent to Eq. (13). Equation (27) can now be rewritten to the form
$\dot{Y}={}-{{A\sqrt{2F_{,R}}}\over{2M_{\mathrm{Pl}}^{3}}}\,\dot{R}^{2}\,.$ (29)
@@
$\phi=\sqrt{{{3}\over{2}}}M_{\mathrm{Pl}}\ln\left[{{2}\over{M^{2}_{\mathrm{Pl}}}}F_{,\sigma}\right]$ (30)
@@
$S_{E}=\int d^{4}x\sqrt{-g^{E}}\left[{{M^{2}_{\mathrm{Pl}}}\over{2}}R_{E}-{{1}\over{2}}g^{E\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-V_{E}(\phi)\right],$ (31)
@@
Varying the action (31) gives the following equations of motion in the spatially flat FLRW universe:
$3\,M^{2}_{PL}\,H_{E}^{2}={{1}\over{2}}\left({{d\phi}\over{dt_{E}}}\right)^{2}+V_{E}\,$, (32)
${{dH_{E}}\over{dt_{E}}}={}-{{1}\over{2M_{PL}^{2}}}\left({{d\phi}\over{dt_{E}}}\right)^{2}\,$, (33)

### v2

# Sergei V. Ketov11footnotetext: The corresponding author.    Ekaterina O. Pozdeeva    Sergey Yu. Vernov

@@ Abstract
The $F(R)$ gravity models of inflation are revisited in light of the recent observations of cosmic microwave background radiation by Atacama Cosmology Telescope (ACT) and DESI Collaboration. A detailed study of the evolution equations in the Jordan frame is given and a new description of the slow-roll approximation in the $F(R)$-gravity-based models of inflation is proposed. It is found that all those models of inflation are significantly constrained by demanding a higher (than the Planck Telescope value) cosmological tilt $n_{s}$ of scalar perturbations and a positive running index $\alpha_{s}$ favored by ACT. It is not difficult to meet the ACT constraints on the scalar tilt $n_{s}$ by modifying the existing models of inflation, but simultaneously demanding a positive running $\alpha_{s}$ would rule out many of them. Using the proposed slow-roll approximation in the Jordan frame, we provide a new modification of the Starobinsky inflation model in the framework of $F(R)$ gravity, which satisfies all ACT constraints. An extension of our ACT-consistent inflation model to the unified $F(R)$-gravity description of Starobinsky-like inflation and production of primordial black holes on a smaller scale is also proposed.
@@ 1 Introduction
Cosmological inflation is a paradigm that proposes the existence of a quasi-de-Sitter phase in the early Universe before the radiation-dominated era (approximately between $10^{-36}$ s to $10^{-32}$ s). Inflation is characterised by an accelerating expansion of the Universe with a ”Graceful Exit”, and resolves major internal problems in the standard (Einstein-Friedmann) cosmology, such as the flatness problem, the horizon problem, the initial conditions problem, etc. The driver of inflation is supposed to be a scalar field called inflaton whose microscopic quantum fluctuations become macroscopic (or classical) during inflation and seed the large-scale structure formation in the Universe. A success of inflation is not limited to the above because it also predicts a nearly scale-invariant and Gaussian spectrum of primordial perturbations in very good agreement with precision observations of the cosmic microwave background (CMB) radiation [1, 2]. Inflation is still a small window into the very early Universe and very high energy physics, while the detailed mechanism of inflation is unknown despite many (over 300) theoretical models proposed in the literature [3].
Single-field models of inflation can provide a simple description of viable inflation in terms of canonical inflaton field $\phi$ minimally coupled to gravity and having a potential $V(\phi)$. The variety of single-field inflation models is thus parameterised by their potentials, whereas CMB observations still allow infinitely many possible potentials, even though the latter have the vanishing measure in the space of all potentials. Therefore, data-driven selection of inflaton potentials needs to be complemented by fundamental theoretical principles and constraints to discriminate between the CMB-compatible potentials.
A scalar field theory minimally coupled to Einstein gravity (called the Einstein frame) is known to be classically equivalent to modified $F(R)$ gravity (called the Jordan frame) whose Lagrangian is given by a smooth function $F(R)$ of the spacetime Ricci curvature scalar $R$, see e.g. Ref. [4]. Both descriptions are related by a Weyl transformation of spacetime metric. The relation between those frames is non-trivial, and it can be used for imposing restrictions on the inflaton potential from the gravity side. Moreover, a description of inflation in pure $F(R)$ gravity without extra matter uses only gravitational interactions with general covariance at work, while it can be technically more advantageous by using only analytical methods instead of numerical calculations. The famous example is the Starobinsky inflation model defined in the Jordan frame by the action [5, 6]
$S_{\rm Star.}={{M^{2}_{\rm Pl}}\over{2}}\int d^{4}x\sqrt{-g}\left(R+{{1}\over{6m^{2}}}R^{2}\right)~,$ (1.1)
@@
$V_{\rm Star.}(\phi)={{3}\over{4}}M^{2}_{\rm Pl}m^{2}\left[1-\exp\left(-\sqrt{\hbox{\large{${{{{2}}\over{{3}}}}$}}}\phi/M_{\rm Pl}\right)\right]^{2}$ (1.2)
that nicely fits the CMB observations by Planck/BICEP [1, 2], though the origin of the potential (1.2) without the reference to the action (1.1) becomes obscure. The Starobinsky model in the Jordan frame is not only geometrical by relying on General Relativity but also explains the origin of the approximate flatness of the potential $V(\phi)$ during inflation due to the scale invariance of the $R^{2}$-action [7], and has an attractor solution for inflation too, which sharply distinguishes the potential (1.2) among others.
At the same time, the Starobinsky description of inflation is expected to provide merely the leading (or dominant) terms in the gravitational effective action and in the inflaton potential, respectively, because both of them are subject to corrections from the higher-order terms that are more difficult to calculate. On the one hand, those corrections are important to account for more precise CMB measurements in the future. On the other hand, a more experimental precision requires a more precise theoretical framework for the CMB description. As a result of all that, single-field models of inflation become more and more constrained. The $F(R)$ gravity models of inflation, generalizing the Starobinsky model, were well studied in the literature, see e.g., Refs. [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. More recently, interest to those inflation models renewed due to the new data from the Atacama Cosmology Telescope (ACT) [21, 22] and DESI 2024 [23].
The single-field inflation models become even more constrained when one adds a production of primordial black holes (PBH) at smaller scales [24, 25, 26]. The standard mechanism of PBH production is via engineering a nearly-inflection point of the potential $V$ [27, 28, 29, 30, 31]. However, less is known about PBH production with inflation in the Jordan frame, which may provide additional insights. Quantum loop corrections might invalidate many single-field models of inflation with PBH production [32], while loop calculations rely on a detailed shape of the potential $V$ [33] or that of the $F$-function [34, 35].
@@
Our paper is organised as follows. In section 2, a detailed study of the evolution equations in a generic $F(R)$ gravity theory is given. The correspondence between the Jordan and Einstein frames is systematically described in section 3. The inflation parameters are introduced and computed in sections 4 and 5. Also in section 5, we propose new explicit formulae for the slow-roll approximation in the Jordan frame. The Starobinsky model of inflation and its predictions for the CMB data are reproduced in the Jordan frame in section 6 from the viewpoint of the proposed SR approximation. The CMB measurements by Planck/BICEP Collaborations [2] and Atacama Cosmology Telescope (ACT) [21, 22] combined with DESI 2024 data [23] are discussed in section 7. A new $F(R)$ gravity model of Starobinsky inflation, extending the modified gravity Lagrangian in Eq. (1.1) to a higher-order polynomial with respect to the spacetime scalar (Ricci) curvature and satisfying the ACT/DESI constraints within $1\sigma$, is presented in section 8. We extend our model to the unified description of inflation, PBHs production and dark energy (DE) in section 9. We summarise our conclusions in section 10.
@@ 2 Evolution equations in $F(R)$ gravity models
Let us consider a generic $F(R)$ gravity model in a curved four-dimensional spacetime with the action (in the Jordan frame)
$S_{F}=\int d^{4}x\sqrt{-g}F(R),$ (2.1)
@@
A variation of the action (2.1) with respect to the metric $g_{\mu\nu}$ gives the following equations of motion:
$F_{,R}R^{\mu}_{\nu}-{{1}\over{2}}F\delta^{\mu}_{\nu}-\left(g^{\mu\rho}D_{\rho}\partial_{\nu}-\delta^{\mu}_{\nu}\Box\right)F_{,R}=0\,$, (2.2)
@@
Equation (2.2) leads to the trace equation
$F_{,R}R-2F+3\Box F_{,R}=0\,.$ (2.3)
@@
$3\left(\dot{H}+H^{2}\right)F_{,R}-{{1}\over{2}}F=3HF_{,RR}\dot{R}\,$, (2.7)
while the trace equation (2.3) is a consequence of Eq. (2.7).
To describe cosmic evolution during inflation, we use the e-folds number $N=-\ln(a/a_{e})$, where $a_{e}$ is a constant. Using the relation ${{d}\over{dt}}=-H{{d}\over{dN}}$, we find that Eq. (2.7) is equivalent to
$\left(H^{2}\right)^{\prime}=4H^{2}-{{R}\over{3}}\,,\qquad{R\,}^{\prime}={}-{{\left(R-6H^{2}\right)F_{,R}-F}\over{6H^{2}F_{,RR}}},$ (2.8)
@@
It is possible to rewrite Eq. (2.7) to the form similar to the $(0,0)$-equation in the Einstein frame, by using the procedure proposed in Ref. [37] and the following functions:
$Y={{M_{\mathrm{Pl}}}\over{\sqrt{2F_{,R}}}}\left(H+{{F_{,RR}\dot{R}}\over{2F_{,R}}}\right),$ (2.9)
$A={{3M_{\mathrm{Pl}}^{4}F_{,RR}^{2}}\over{4F_{,R}^{3}}}\,,\qquad V_{\mathrm{eff}}={{M_{\mathrm{Pl}}^{4}}\over{4F_{,R}^{2}}}\left(RF_{,R}-F\right)\,.$ (2.10)
Equation (2.7) implies that these functions satisfy the equation
$3M_{\mathrm{Pl}}^{2}Y^{2}={{A}\over{2}}\dot{R}^{2}+V_{\mathrm{eff}}$ (2.11)
@@
$\begin{split}3M_{\mathrm{Pl}}^{2}Y^{2}=&{{3M_{\mathrm{Pl}}^{4}}\over{2F_{,R}}}\left(H^{2}+{{HF_{,RR}\dot{R}}\over{F_{,R}}}+{{F_{,RR}^{2}}\over{4F^{2}_{,R}}}\dot{R}^{2}\right)\\ =&{{3M_{\mathrm{Pl}}^{4}}\over{2F_{,R}}}H^{2}+{{3M_{\mathrm{Pl}}^{4}}\over{2F_{,R}^{2}}}\left[\left(\dot{H}+H^{2}\right)F_{,R}-{{F}\over{6}}\right]+{{A}\over{2}}\dot{R}^{2}\\ =&{{A}\over{2}}\dot{R}^{2}+V_{\mathrm{eff}}\,.\end{split}$ (2.12)
@@ 3 Correspondence between Jordan and Einstein frames
Let us demonstrate that the function $Y(t)$ is the Hubble parameter in the Einstein frame as a function of the cosmic time in the Jordan frame. The $F(R)$ gravity action (2.1) can be rewritten as
$S_{J}=\int d^{4}x\sqrt{-g}\left[U(\sigma)R-V(\sigma)\right]\,$, (3.1)
@@
$U(\sigma)=F_{,\sigma}(\sigma),\qquad V(\sigma)=\sigma F_{,\sigma}-F\,$, (3.2)
and the commas denote the derivatives with respect to a given scalar field. When $F_{,\sigma\sigma}(\sigma)\neq 0$, the field $\sigma$ can be eliminated via its algebraic equation of motion, $R=\sigma$, which yields back the action (2.1). It follows from the definitions of $U$ and $V$ that
$V_{,\sigma}=\sigma U_{,\sigma}=\sigma F_{,\sigma\sigma}\,.$ (3.3)
@@
$2U\dot{H}-H\dot{U}+\ddot{U}=0.$ (3.5)
After the Weyl transformation of the metric from the Jordan frame to the Einstein frame as
$g^{(E)}_{\mu\nu}={{2F_{,\sigma}(\sigma)}\over{M^{2}_{\mathrm{Pl}}}}g_{\mu\nu}\,$, (3.6)
@@
$S_{E}=\int d^{4}x\sqrt{-g^{(E)}}\left[{{M^{2}_{\mathrm{Pl}}}\over{2}}R_{E}-{{h}\over{2}}{g^{(E)\mu\nu}}\partial_{\mu}{\sigma}\partial_{\nu}{\sigma}-V_{eff}\right],$ (3.7)
@@
$h(\sigma)={{3M^{2}_{\mathrm{Pl}}}\over{2F_{,\sigma}^{2}}}F_{,\sigma\sigma}^{2}={{2U(\sigma)}\over{M^{2}_{\mathrm{Pl}}}}\,A(\sigma)$ (3.8)
@@
$V_{eff}(\sigma)={{M_{\mathrm{Pl}}^{4}}\over{4F_{,\sigma}^{2}}}\left(RF_{,\sigma}-F\right)={{M_{\mathrm{Pl}}^{4}V}\over{4U^{2}}}~.$ (3.9)
@@
$H_{E}={{d\ln(a_{E})}\over{dt_{E}}}={{M_{\mathrm{Pl}}}\over{\sqrt{2U}}}\left(H+{{\dot{U}}\over{2U}}\right),$ (3.10)
@@
${{dt_{E}}\over{dt}}={{\sqrt{2F_{,R}}}\over{M_{\mathrm{Pl}}}}\,.$ (3.11)
Equations (3.4) and (3.5) can be rewritten in terms of $H_{E}$ as
$3M_{\mathrm{Pl}}^{2}H_{E}^{2}={{A}\over{2}}{\dot{\sigma}}^{2}+V_{\mathrm{eff}},$ (3.12)
$\dot{H}_{E}={}-{{A\sqrt{2U}}\over{2M_{\mathrm{Pl}}^{3}}}\,{\dot{\sigma}}^{2},$ (3.13)
where the functions $A(\sigma)$ and $V_{eff}(\sigma)$ have been defined in Eq. (2.10).
Differentiating Eq. (3.12) with respect to the time $t$ and using Eq. (3.13), one yields
$\ddot{\sigma}={}-3\sqrt{{{2U}\over{M_{\mathrm{Pl}}^{2}}}}H_{E}\dot{\sigma}-{{A_{,\sigma}}\over{2A}}{\dot{\sigma}}^{2}-{{V_{\mathrm{eff},\sigma}}\over{A}}.$ (3.14)
Using Eqs. (3.2) and (3.10), one gets $H_{E}=Y$, so that Eq. (3.12) is equivalent to Eq. (2.11). Equation (3.13) can now be rewritten to the form
$\dot{Y}={}-{{A\sqrt{2F_{,R}}}\over{2M_{\mathrm{Pl}}^{3}}}\,\dot{R}^{2}\,.$ (3.15)
@@
$\phi=\sqrt{{{3}\over{2}}}M_{\mathrm{Pl}}\ln\left[{{2}\over{M^{2}_{\mathrm{Pl}}}}F_{,\sigma}\right]$ (3.16)
@@
$S_{E}=\int d^{4}x\sqrt{-g^{E}}\left[{{M^{2}_{\mathrm{Pl}}}\over{2}}R_{E}-{{1}\over{2}}g^{E\mu\nu}\partial_{\mu}\phi\partial_{\nu}\phi-V_{E}(\phi)\right],$ (3.17)
@@
A variation of the action (3.17) gives the following equations of motion in the spatially flat FLRW universe:
$3M^{2}_{\mathrm{Pl}}H_{E}^{2}={{1}\over{2}}\left({{d\phi}\over{dt_{E}}}\right)^{2}+V_{E}\,$, (3.18)
${{dH_{E}}\over{dt_{E}}}={}-{{1}\over{2M_{\mathrm{Pl}}^{2}}}\left({{d\phi}\over{dt_{E}}}\right)^{2}\,$, (3.19)